# A cross-platform analysis of polarization and echo chambers in climate change discussions

### v2

@@ II Twitter analysis
To evaluate the isolation of the communities, we calculated the directed interactions between them, and we computed the normalised outflow difference given by
$P^{out}_{i}=\frac{\sum_{j,j\neq i}T_{ij}-T_{ii}}{\sum_{\forall j}T_{ij}},$ (1)
and the entropy of flows as
@@
where $T_{ij}$ is the number of interactions from community $i$ to community $j$ and $p_{ij}$ is the probability of having a retweet from $i$ to $j$ given by $T_{ij}/\sum_{\forall j}T_{ij}$. While the first index quantifies the ratio between external and internal links and goes from $-1$ to $1$ when all flows are external, the entropy provides information on the variety of those flows. The entropy is 00 when there is no variety of flows and $1$ when it is maximum.
In Fig. 4, we represent the outflow entropy as a function of the isolation for the four networks. Most of the largest communities exhibit a certain degree of isolation, indicated by negative values. Across all datasets, the right-biased community demonstrates strong isolation and low entropy, implying their messages have limited spread across the network. However, the pattern changes when considering the mention graph (Fig. S4). The largest mainstream communities show minimal mentions of other communities, whereas the low-reliability communities display lower isolation and more external mentions. This observation suggests that climate change sceptics may reference the mainstream community to attract attention but get little attention instead.
Figure 4: Community connectivity and entropy in retweet networks. Outflow entropy as a function of the normalised outgoing isolation in the retweet networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of feasible values.
We computed the minimum and maximum attainable entropy values as a function of $P^{out}_{i}$ and $p_{ii}$ for orientation. Given a value of $P^{out}_{i}$ we have
$\frac{\sum_{j,j\neq i}T_{ij}}{{\sum_{\forall j}}T_{ji}}=P_{i}+\frac{T_{ii}}{{\sum_{\forall j}T_{ij}}}=P_{i}+p_{ii}.$ (3)
The minimum value of the entropy corresponds to the situation when a community is connected only to itself and one external community (except for the extreme case $P_{i}=1$), which we can plug into the entropy as
$H^{in}_{i}=-p_{ii}logp_{ii}-P_{i}+p_{ii}logP_{i}+p_{ii}.$ (4)
$H^{in}_{i}=-p_{ii}logp_{ii}-\sum^{N_{c}-1}_{j=0}\frac{P_{i}+p_{ii}}{N_{c}-1}log\frac{P_{i}+p_{ii}}{N_{c}-1}.$ (5)
In Fig. S5, we present the reversed quantities computed for the inflow, where we also observe that the right-biased communities with low reliability are among the most isolated.
@@
Given that many networks include users from different countries speaking various languages, we also calculated polarization by focusing on the two largest communities speaking the same language with opposing biases, comparing the networks of retweets and mentions (Fig. S6). They still exhibit polarization in the retweet network, however, the scores are considerably smaller or even negative in the mention network. Thus, we confirm the previous findings on the disparity in network organization based on interaction typology. Our results suggest that focusing solely on retweet networks may provide a biased perspective on Twitter dynamics and that effective polarization, in terms of interaction between opposing views, may be low.
II.1 User analysis
@@
Throughout this section, we assess the presence of echo chambers with a focus on the two largest communities exhibiting opposing political biases. We employed an approach developed in [31] to evaluate the existence of echo chambers. For a given user $i$ with degree $k_{i}$, we computed its average bias $x_{i}$ and reliability $y_{i}$, along with the average bias and reliability of its neighbourhood, denoted as $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}x_{j}$ and $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}y_{j}$, respectively.
@@
Figure 5: Detection of bias echo chambers in the retweet networks. Neighbourhood political bias as a function of the user bias in the retweet network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
@@
In Fig. 5, we show the density plot for the bias of the neighbourhood of users as a function of their own values when considering the interactions through retweets. As expected from the overall ideological positioning of the users, there is a hotspot on the left of the ideological spectrum and a weaker one on the right side. We have identified a diagonal trend that could be related to the presence of echo chambers. Right-biased users interact more with left-biased users, as there are two hotspots for positive values in some cases. The presence of echo chambers is more evident in the Doñana dataset than in the rest of the networks, mainly due to the presence of more users with right bias. There are three hotspots of comparable magnitude across the diagonal with right, centre, and left bias. Still, the hotspot on the left side of the spectrum displays a lower ideological isolation, since it interacts with users featuring a bias close to 0. The right-biased users interact mainly between themselves and not with users in the centre. When analysing other types of interactions, such as the mentions (Fig. S7), the presence of echo chambers is weaker with nearly horizontal hotspots in the COP27 and IPCC datasets. This observation is consistent with the results on polarization from the previous section. The existence of echo chambers is less pronounced when considering different types of interactions, and mentions could serve as a bridge between users with different political biases.
@@
We also examine the echo chambers of the content reliability in Fig. S8, where there are three hotspots for high, medium, and small values. The reliability has a similar behaviour, the hotspots in high reliability suggest that users who share highly reliable sources interact with each other, while low-reliability users interact with users that spread highly reliable sources.
Figure 6: Chamber overlap distribution $q_{ij}$. Chamber overlap distribution for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana calculated on the top $20$ users of each community.
@@
We inspect in detail the user interaction similarity by quantifying their echo chamber [81], without considering the ideology. Given a set of leading users $i$, which have an audience $A_{i}$ composed of the set of users that retweeted it, their chamber $C_{i}$ is the set of users retweeted by audience $A_{i}$. We focus on the two communities with opposing political biases, specifically the top 20 users with the most retweets in each of those communities. The first quantity we will analyse is the chamber overlap between a user $i$ and a user $j$ given by
@@
where $C_{i}$ and $C_{j}$ are the chambers of users $i$ and $j$. It is worth noting that the quantity $q_{ij}$ is symmetric by definition and is equal to $1$ when chambers are identical and 00 when they are completely different.
@@
In Fig. 6, we show the chamber overlap distribution $P(q_{ij})$, where a notable disparity between networks is observed. A strong bimodal distribution suggests the presence of echo chambers, as each user has low values of $q_{ij}$ with users outside its echo chamber and high values with those inside it. The results from the COP27 and Climate Refugees datasets suggest that leaders have more heterogeneous chambers compared to the IPCC and Doñana datasets, where we observe a peak for high chamber overlap. The Doñana dataset features a strong echo chamber effect with two separated peaks at low and high values of $q_{ij}$. In Fig. S9, we have performed hierarchical clustering of the top 20 users per community based on the similarity of their chambers, showing a strong alignment between the communities detected.
II.2 Hashtag analysis
Next, we inspect the patterns of hashtags shared in each network, focusing on the two largest communities with opposing biases. We report in Fig. S10 the top hashtags per network, where we show that despite the ideological confrontation, they share a large number of them. The mainstream communities use rather generalistic hashtags such as #$\#COP27$, #$\#IPCC$, #$\#climatechange$, #$\#ar6$, or #$\#climatereport$. The results suggest that the users of the right-biased community use common hashtags to attract the attention of mainstream users. In the case of the Doñana, we observe hashtags related to the environment as #$\#donana$, #$\#donanaseextingue$, or #$\#salvemosdonana$ for the mainstream community. The community with right bias has hashtags related to right-wing parties as #$\#teamvox$, or against the left-wing government as #$\#gobiernodimision$. Common hashtags between communities are less frequent in the Doñana dataset in line with the more pronounced echo chambers detected.
To compute the hashtag polarization, we calculated the fraction of tweets with each hashtag by community given by
$P_{i}^{\#}=\frac{N_{i}^{r}-N_{i}^{l}}{N^{l}_{i}+N_{i}^{r}},$ (7)
where $N_{i}^{l}$ and $N_{i}^{r}$ are the number of tweets with the hashtag $i$ in the main left and right-wing communities, respectively. We display the polarization of the top hashtags in Fig. 7. The number of hashtags in each network varies since we have considered the $20$ most shared in each community, and some of them can overlap. Interestingly, we observe only a few hashtags with values close to 00, suggesting a strong polarization in the use of hashtags. In the IPCC and COP27 datasets, the mainstream community dominates most of the hashtags, while the community spreading a counter-narrative dominates only a few hashtags related to the denial of climate change that include #$\#climatehoax$ or #$\#agenda2030$. In COP27, we also observe hashtags related to hypocrisy as #$\#greenwash$. In the Doñana network, we also observe only a few hashtags close to zero and hashtags almost equally divided between $+1$ and $-1$. We have analysed the chamber overlap of hashtag dynamics in Fig. S11.
@@ II.2 Hashtag analysis
Figure 7: Hashtag polarization. Polarization around hashtags for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. Negative and positive values correspond to hashtags more abundant in the left-biased and right-biased communities, respectively.
III Youtube analysis

### v3

@@ II.1.1 Isolation and entropy
We further evaluate the connectivity patterns between communities by quantifying the structural properties of their inter- and intra-group interactions. To quantify the isolation of communities, we calculated the directed interactions between them, and we computed the normalised outflow difference given by
$\Delta F^{out}_{i}=\frac{\sum_{j,j\neq i}T_{ij}-T_{ii}}{\sum_{\forall j}T_{ij}},$ (1)
where $T_{ij}$ is the number of interactions from community $i$ to community $j$. $\Delta F^{out}_{i}$ provides a measure of a community’s isolation. A value of 1-1 indicates that a community has only internal flow, whereas a value close to $1$ suggests that most of the flow is directed toward other communities. A value of 0 signifies that half of the flow is external to the community, while the other half remains internal.
To understand the diversity of interactions among communities, we compute the entropy of flows as
@@
where $p_{ij}$ is the probability of having an interaction from $i$ to $j$ given by $T_{ij}/\sum_{\forall j}T_{ij}$. While $\Delta F^{out}_{i}$ quantifies the ratio between external and internal links and goes from 1-1 (when all flows are internal) to $1$ (when all flows are external), the entropy ($H^{out}_{i}$) provides information on the variety of those flows. The entropy is 0 when there is no variety of flows and $1$ when it is maximum. Most of the largest communities exhibit a certain degree of isolation, indicated by negative values. In the COP27 network, the most isolated communities are 2 and 3, both Brazilian ($\Delta F^{out}_{2}=-0.97$ and $\Delta F^{out}_{3}=-0.94$), followed by community 1 ($\Delta F^{out}_{1}=-0.88$), which is the most isolated and has the lowest entropy ($H^{out}_{1}=0.05$) among English-speaking groups. This community consists of sceptics with influential accounts such as @JamesMelville and @BernieSpofforth. Following the same pattern, in the IPCC network, community 1 has the highest isolation value ($\Delta F^{out}_{1}=-0.98$) and the lowest entropy ($H^{out}_{1}=0.014$). This community also includes, among other accounts, @BernieSpofforth. Similarly, in the Climate Refugees network, community 6 is the most isolated ($\Delta F^{out}_{6}=-0.93$) and has the lowest entropy ($H^{out}_{6}=0.03$), featuring accounts such as @MaximeBernier and @PrisonPlanet. In the Spanish Doñana network, community 0 is the most isolated ($\Delta F^{out}_{0}=-0.95$) and has the lowest entropy ($H^{out}_{0}=0.02$). Across all datasets, communities with right-wing alignment demonstrate strong isolation and low entropy, implying that their messages have limited spread across the network.
However, the pattern changes when considering the mention graph. The largest mainstream communities show far less mentions of other communities, for instance, community 0 on IPCC and COP27 networks has high isolation ($\Delta F^{out}_{0}=-0.5$ and $\Delta F^{out}_{0}=-0.57$, respectively). Whereas the low-reliability communities display lower isolation and more external mentions, its the case of community 1 in COP27 ($\Delta F^{out}_{1}=0.37$ and $H^{out}_{1}=0.25$). These results reinforce, first, the heterogeneity in group connectivity depending on the interaction types and, second, different levels of isolation by communities.
II.2 Political polarisation
II.2.1 Bias and reliability
We assess if the structural polarisation is related to the divergence in political views and information diet by calculating the bias and reliability of the URLs shared by the users. Denser connections between users with similar ideological biases could indicate a closer influence between them and a greater disagreement on climate change discussions. Information on the political bias and reliability of news outlets was obtained from MediaBias fact-check [74], where positive and negative political bias correspond to right-wing and left-wing media, respectively. Outlets with higher reliability are considered more trustworthy and tend to share more verified news. For the Doñana dataset, the reliability and bias were based on a separate dataset [75], given that most media sources are Spanish. To facilitate the analysis, we have normalised the bias within the range [-1, 1] and the reliability within the range [0, 1]. The minimum and maximum biases are set according to the higher absolute value observed. In Fig. 3(a)-(d), we show the density plot of community reliability as a function of their political bias according to the shared URLs in the IPCC, COP27, Climate Refugees, and Doñana networks. We computed user-level values by averaging over the sources they shared and then calculated community-level values by averaging over the users. All datasets exhibit a left-wing bias, as most communities show negative values, especially in the IPCC and Climate Refugee networks. Possible explanations include a systemic bias in media coverage of these topics, the tendency of climate change sceptics to share information from unconventional sources that are not explicitly flagged by media bias indicators, or an over-representation of left-wing users in climate change discussions, as they may be more likely to engage due to greater awareness of the issue. The left-wing bias is particularly pronounced in the IPCC and Climate Refugee networks, where the cluster of large communities is centred around negative values. However, the bias is more dispersed in the Doñana dataset, possibly due to differences in the shared media sources or a wider ideological variety.
The communities with right-wing bias we have identified are: community 1 in COP27 (with key actors as @JamesMelville) and IPCC (with the presence of @BernieSpofforth), community 6 in Climate Refugees (with @MaximeBernier or @PrisonPlanet), and community 0 in Doñana (with @JuanMa_Moreno or @alfonso_ussia). We performed significance tests to compare the bias and reliability of communities, revealing several significant differences. In terms of bias, the most notable differences (p-value ¡ 0.001) are observed in the previously mentioned right-wing biased communities compared to the other communities. Regarding reliability, community 4 in COP27 (with main actors such as @CarolineLucas), community 3 in IPCC (with @GretaThunberg or @CarolineLucas), community 5 in Climate Refugees (@jeremycorbyn), and community 1 in Doñana (@WWFespana and other conservation organisations) show the most significant higher reliability relative to the other communities (see Figs. S6-S7).
Based on the previous findings and manual inspection of the communities, we identified the two largest communities with opposing political leanings. In the IPCC and COP27 datasets, community 0 represents the largest community aligned with the mainstream climate change narrative, while community 1 is the primary advocate of a counter-narrative message. However, in the Doñana network, the labels are swapped, as the right-wing biased community has a larger user base. For the Climate Refugees network, most of the top 10 communities exhibit similar bias-reliability values, except for community 6, which stands out with low reliability and a right-wing bias. In this dataset, we designate community 1 as the primary left-wing biased community, as it includes well-known climate change activists such as @MikeHudema or @GretaThunberg and engages in more interactions with the low-reliability community. Our results point out that most communities in climate change discussions share reliable information and have a left-biased alignment. However, there is at least one large community in each dataset on the right-side alignment that shares less reliable information.
These results indicate differences in the network topology depending on the type of interactions. Retweets, which are linked to influence, connect individuals in communities with similar biases. Instead, a higher number of mentions across communities suggests interactions between individuals from opposing sides of the ideological spectrum.
@@ II.2.1 Bias and reliability
Given that many networks include users from different countries speaking various languages, we also measured polarisation by focusing on the two largest communities that speak the same language but have opposing biases, comparing the retweet and mention networks. The results still indicate polarisation in the retweet network, with modularity values ranging from $0.07$ in the IPCC network to $0.21$ in the Doñana network. Even in the mention networks—where structural scores are lower (modularity is $0.05$ in COP27, $0.16$ in Climate Refugees, and $0.15$ in Doñana)—only the IPCC network shows no polarisation (modularity = 0.23-0.23). These findings further support the observed disparity in network organisation based on interaction typology (see Section S2.3 for more details).
II.2.2 Echo chambers
@@
Throughout this section, we assess the presence of echo chambers with a focus on the two largest communities exhibiting opposing political biases. Our objective is to assess the interaction patterns between users as a function of their bias. We employed an approach developed in [9] to evaluate the existence of echo chambers. For a given user $i$ with degree $k_{i}$, we computed its average bias $x_{i}$ and reliability $y_{i}$, along with the average bias and reliability of its neighbourhood, denoted as $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}x_{j}$ and $\frac{1}{k_{i}}\sum_{\forall j\neq i}A_{ij}y_{j}$, respectively.
@@
Figure 4: Detection of echo chambers in ideological biases for the retweet networks. Neighbourhood political bias as a function of the user bias in the retweet network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
@@
We show the density plot for the bias of the neighbourhood of users as a function of their values when considering the interactions through retweets in Fig.4. As expected from the overall ideological positioning of the users, there is a hotspot on the left of the ideological spectrum and a weaker one on the right side. We have identified a diagonal trend that could be related to the presence of echo chambers. Right-wing biased users interact more with left-wing biased users, as there are two hotspots for positive values in some cases. The presence of echo chambers is more evident in the Doñana dataset than in the rest of the networks, mainly due to the presence of more users with right-leaning discourse. There are three hotspots of comparable magnitude across the diagonal with right-, centre-, and left-leaning discourses. Still, the hotspot on the left side of the spectrum displays a lower ideological isolation, since it interacts with users featuring a bias close to 0. The right-wing biased users interact mainly with each other and not with users in the centre. When analysing other types of interactions, such as the mentions, the presence of echo chambers is weaker with nearly horizontal hotspots in the COP27 and IPCC datasets (see details in Figs. S11). This observation is consistent with the results on polarisation from the previous section. The existence of echo chambers is less pronounced when considering different types of interactions, and mentions could serve as a bridge between users with different political biases.
@@
We also examine the echo chambers based on the reliability of the content shared by users and their neighbourhoods, where we find three hotspots for high, medium, and small values. The hotspots in high reliability suggest that users who share highly reliable sources interact with each other, while low-reliability users interact with users who spread highly reliable sources (see Fig. S12).
@@
We inspect in detail the user interaction similarity by quantifying their echo chamber [76], without considering the ideology. Given a set of leading users $i$, which have an audience $A_{i}$ composed of the set of users that retweeted it, their chamber $C_{i}$ is the set of users retweeted by audience $A_{i}$. We focus on the two communities with opposing political biases, specifically the top twenty users with the most retweets in each of those communities. The first quantity we will analyse is the chamber overlap between a user $i$ and a user $j$ given by
@@
where $C_{i}$ and $C_{j}$ are the chambers of users $i$ and $j$. It is worth noting that the quantity $q_{ij}$ is symmetric by definition and is equal to $1$ when chambers are identical and 0 when they are completely different.
@@
We compute the chamber overlap distribution $P(q_{ij})$, where a notable disparity between networks is observed (see Fig.S13). A strong bimodal distribution suggests the presence of echo chambers, as each user has low values of $q_{ij}$ with users outside its echo chamber and high values with those inside it. The results from the COP27 and Climate Refugees datasets suggest that leaders have more heterogeneous chambers compared to the IPCC and Doñana datasets, where we observe a peak for high chamber overlap. The Doñana dataset features a strong echo chamber effect with two separated peaks at low and high values of $q_{ij}$. We have performed hierarchical clustering of the top 20 users per community based on the similarity of their chambers, showing a strong alignment between the communities detected (for more details see Section S2.6).
@@
Figure 5: Hashtag polarisation. Polarisation around hashtags for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. Negative and positive values correspond to hashtags more abundant in the left-biased and right-biased communities, respectively.
II.3 Policy and affective polarisation
On account of the structural polarisation observed, with groups of users sharing political bias, we further investigate if there are perceptible differences in their climate change views. Structural metrics are purely based on user interactions and, thus, provide limited information on the reasons and motivations of the groups. To get a first assessment of the content discussed in each of the two opposed communities and evaluate if they have confronted perspectives on climate change, we inspect the patterns of hashtags shared. Despite their ideological confrontation, they share most of them, likely due to the common subject of discussion (Fig. S15). The mainstream communities use rather generalist hashtags such as #COP27\#COP27, #IPCC\#IPCC, #climatechange\#climatechange, #ar6\#ar6, or #climatereport\#climatereport. In Doñana, common hashtags between communities are less frequent, which could indicate a stronger political polarisation. We observe hashtags related to the environment as #donana\#donana, #donanaseextingue\#donanaseextingue, or #salvemosdonana\#salvemosdonana for the mainstream community and hashtags related to right-wing parties as #teamvox\#teamvox, or against the left-wing government as #gobiernodimision\#gobiernodimision, in the right-wing biased community.
To compute the hashtag polarisation, we calculated the fraction of tweets with each hashtag by community, given by
$P_{i}^{\#}=\frac{S^{l}N_{i}^{r}-S^{r}N_{i}^{l}}{S^{r}N^{l}_{i}+S^{l}N_{i}^{r}},$ (4)
where $N_{i}^{l}$ and $N_{i}^{r}$ are the number of tweets with the hashtag $i$ in the main left- and right-wing communities, respectively, and $S^{l}$ and $S^{r}$ the number of users in each community. This formulation takes into account the uneven sizes of the communities. We display the polarisation of the top hashtags in Fig. 5. The number of hashtags in each network varies since we have considered the $20$ most shared in each community, and some of them can overlap. Interestingly, we observe only a few hashtags with values close to 0, suggesting a strong polarisation in the use of hashtags. In the IPCC and COP27 datasets, the mainstream community dominates most of the hashtags, while the community spreading a counter-narrative dominates only a few hashtags related to the denial of climate change that include #climatehoax\#climatehoax or #agenda2030\#agenda2030. In COP27, we also observe hashtags related to environmental hypocrisy such as #greenwash\#greenwash. In the Doñana network, we also observe only a few hashtags close to zero and hashtags almost equally divided between +1+1 and 1-1. We performed a Z-test to assess the significance of hashtag frequency differences between the most polarised communities. Among the hashtags with highest absolute Z-scores, #climatescam\#climatescam is predominantly used by right-leaning communities in both the COP27 and IPCC networks. In the Climate Refugees network, one of the hashtags with a high Z-score is #howdareyou\#howdareyou, again primarily used by the right-wing community. In the case of the Spanish discussion about Doñana, the hashtags with the highest Z-scores are those criticising the Spanish government, as previously mentioned (see Section S3.1 for details).
Whereas hashtags can provide a glimpse of the ideas discussed within each community, they offer a limited perspective to the opinion of individuals as they can be used in different contexts. To better understand the divergence in climate change views, we conducted a topic analysis on tweets from the two largest communities with opposing political perspectives. We assess topic polarisation by quantifying the number of users within each community who discuss a given topic, using Eq. 4. Most topics are either dominated by the left-biased community or the low-reliability community. In the COP27 network, the right-leaning discourse reveals corruption discussions with money-related terms, and political references to the UK, China and India. In contrast, the left-biased community focuses on essential resources such as water and food, as well as the urgency of taking action. In the IPCC network, topics within the right-biased community express scepticism about the scientific basis of climate change predictions, reference COVID-19, and include discussions that are often unrelated or misleading (for more details about the topics see Section S3.2). The results confirm that it could be composed of individuals posting low-reliability information. The left-biased community emphasises the need to reduce emissions and the urgency of taking action, using language similar to that found in COP27 discussions. In the Climate Refugees network, the right-biased community focuses on jobs and border security, whereas the left-biased community highlights the risks posed by natural disasters. The greatest divergence in discussions is observed in the Doñana network. The right-biased community focuses on the political handling of the situation, with an emphasis on the Spanish president, the ruling party (PSOE), and the European Union. In contrast, the mainstream community centres its discussion on the Doñana National Park and its conservation.
Figure 6: Average negative sentiment in messages in community interactions. Average negative sentiment of the messages as a function of the poster community (y-axis) and quoted (x-axis). The results correspond to (a) COP27, (b) IPCC,(c) Climate Refugees, (d) Doñana.
Besides the disparity in the content shared, via hashtags and topics, we want to adress whether those groups are aware of each other and if they are emotionally aligned. Therefore, We evaluate emotional repulsion in relation to political views by analysing the sentiment of quote messages. Specifically, we assess the average negative sentiment based on the community of both the sender and the receiver (Fig. 6). We have implemented a sentiment analysis on the four Twitter datasets using a RoBERTa-based model fine-tuned on around 58 million of tweets [77, 78]. Overall, messages from right-wing biased communities exhibit a higher degree of negative sentiment, particularly when directed at left-wing biased communities (0.48 for COP27, 0.54 for IPCC, and 0.55 for Climate Refugees). The difference between intra- and inter-community messages is most pronounced in the IPCC and Doñana networks. In the IPCC network, the negativity score for right-leaning to left-leaning interactions is 0.54, while for right-leaning to right-leaning interactions, it is 0.46 ($\Delta=0.08$). Similarly, in the Doñana network, right-leaning to left-leaning interactions have a negativity score of 0.38, compared to 0.31 for right-leaning to right-leaning interactions ($\Delta=0.07$). Interestingly, the most pronounced difference in sentiment between intra- and inter-community interactions occurs among users in the left-wing biased community. In the extreme case of the COP27 network, left-leaning users display a $\Delta=0.23$ difference, while also being associated with the dissemination of more reliable information. Across all networks, the lowest levels of negative sentiment are observed within the left-wing biased communities, ranging from 0.20 in COP27 to 0.39 in Climate Refugees. In contrast, the right-wing biased community exhibits consistently higher negativity, even within its own interactions, with values ranging from 0.31 in Doñana to 0.53 in Climate Refugees.
We also conducted a Mann-Whitney U test to evaluate the statistical significance of negativity differences between communities. A consistent pattern emerges in the COP27, IPCC, and Climate Refugees networks: right-leaning to left-leaning interactions display significantly greater negativity than all other combinations. For example, in the IPCC network, the negativity of right-leaning to left-leaning interactions is significantly greater than that of left-wing biased to right-wing biased interactions ($p<.001$), left-wing biased to left-wing biased interactions ($p<.001$), and right-wing biased to right-wing biased interactions ($p<.001$) (see Fig. S19 for additional comparisons). Furthermore, across all datasets, right-wing biased to right-wing biased interactions exhibit significantly higher negativity compared to left-wing biased to left-wing biased interactions (COP27: $p<.001$; IPCC: $p<.001$; Climate Refugees: $p<.001$; and Doñana: $p<.001$). A similar pattern is observed when comparing left-wing biased to right-wing biased interactions against left-wing biased to left-wing biased interactions (COP27: $p<.001$; IPCC: $p<.001$; Climate Refugees: $p<.001$; and Doñana: $p<.001$). Hence, the structural communities identified are aligned with a divergence on content discussion related to climate change and stronger emotional responses. The results suggest that polarisation on the subject takes multiple forms.
II.4 Cross-platform analysis